the ticket at a dance recital costs $5.00 for adults and $2.00 for children. if the total number of tickets sold was 295 and the total amount collected was $1220, how many adults ticket and how many children tickets were sold?

Answer by checkley79(3341)

A+C=295
A=295-C
5A+2C=1220
5(295-C)+2C=1220
1475-5C+2C=1220
-3C=1220-1475
-3C=-255
C=-255/-3
C=85 CHILDREN TICKETS WERE SOLD.
A=295-85=210 ADULT TICKETS WERE SOLD.
PROOF:
5*210+2*85=1220
1050+170=1220
1220=1220